# TwinVLA: Data-Efficient Bimanual Manipulation with Twin Single-Arm Vision-Language-Action Models

### v1

@@ 4 TwinVLA
4.1 Single-Arm Policy Duplication
To construct TwinVLA from SingleVLA, we initialize the twin policies for the left and right arms by copying the pretrained SingleVLA. However, instead of duplicating the full model, we share the vision encoder and DiT (Peebles2022ScalableDM) action head while fully replicating the VLM. Each arm has its own lightweight proprioception encoder. This design yields a compact 1.31.3B-parameter model, comparable to the 1.21.2B-parameter RDT-1B, without significantly increasing computational cost.
4.2 Joint Attention for Cross-arm Fusion
We integrate arm-specific inputs using a joint attention mechanism inspired by Mixture of Transformers (MoT) (liang2024mixtureoftransformerssparsescalablearchitecture). As illustrated in Figure˜3(b), this is achieved by sharing only the self-attention layers across the VLM backbones, while other components like the projection networks operate independently for each arm. Unlike $\pi_{0}$ (black2024pi_0), which links VLM with an action head, we connect two VLMs directly. The detailed pseudocode of joint attention is presented in Algorithm˜1.
@@ 4.3 Mixture-of-Experts Integration
In TwinVLA, each VLM receives both shared and arm-specific inputs to produce single-arm actions. Feeding the shared modality $(l,I_{\text{ego}})_{t}$ redundantly to both VLMs is inefficient, as it increases token length and VRAM usage in a way that hinders the practical training of large models. To address this, we apply a Mixture-of-Experts (MoE) mechanism that routes the shared modality tokens between both VLMs, reducing token length while preserving representational power.
While the MoE effectively shares the feedforward networks (FFNs) between the twin VLM backbones, other key components like projection layers remain duplicated. To address this, we implement an effective form of task arithmetic (10.5555/3692070.3694018). We process an input through the corresponding component in each VLM and then average their outputs. This output-averaging technique functionally simulates a single shared layer without averaging the parameter itself. This is illustrated in the center of Figure˜3(b). Thanks to this change, we enable training the model with a batch size of $8$ on a single 4040GB GPU. Detailed implementation is provided in Appendix C.3.
Attention re-weighting. A potential side effect of introducing new arm-specific tokens is that the model’s learned attention patterns can be disrupted, shifting focus away from the pretrained shared modalities. To mitigate this and preserve the valuable pretrained knowledge, we re-scale the attention scores for the shared modality. This maintains pretrained modality importance, allowing the model to bypass an initial adaptation phase and focus directly on the target task—a benefit evidenced by a lower initial loss and converged loss during fine-tuning.
(a) Real-world tasks
(b) Simulation tasks
Figure 4: Experimental setups. (a) We evaluate TwinVLA on three real-world bimanual tasks using an Anubis robot. (b) We further analyze TwinVLA on a large suite of simulation tasks: 55 tasks in Tabletop-Sim and $50$ tasks in RoboTwin 2.0.
@@ 5 Experiments
  • Does TwinVLA retain core VLA properties—language-following and robustness to unseen scenes and instructions (Sections˜5.5 and 5.6)?
@@ 5.1 Compared Methods
  • RDT-1B (liu2024rdt): This serves as our direct baseline. With a comparable size(1.21.2B vs. TwinVLA’s 1.31.3B parameters), it represents the standard monolithic approach that requires substantially larger resources (1.41.4M trajectories, $\sim$11,$440$ H100 days vs. 0.50.5M single-arm data, $\sim$$25$ H100 days).
  • $\bm{\pi_{0}}$ (black2024pi_0): We include this as a skyline, as this is 3.33.3B-parameter VLA trained on over 1010K hours of proprietary robot data. Our goal is to assess how closely TwinVLA can approach this performance ceiling with far greater efficiency.
  • Diffusion Policy (DP) (chi2023diffusion): This is a strong baseline method in low-data regime with 271271M parameters, used to demonstrate the crucial benefits of pretraining.
@@ 5.2 Real-World Experiments
Environment. For real-world experiments, we use a dual-arm robot, Anubis (anubis2025), as shown in Figure˜4(a). Anubis has two $6$ DoF arms with parallel-jaw grippers. The robot is equipped with two wrist-mounted cameras and a single ego-centric view camera.
Tasks. We design three long-horizon tabletop manipulation tasks which requires careful coordination and accurate motions: carrot to bag, brush to dustpan, and take towel off. We collect $50$ episodes for each task using absolute EEF control. We fine-tune all methods for each task, and evaluate them with $20$ rollouts per task.
Results. As presented in Figure˜5, our model, TwinVLA, significantly outperforms its direct competitor, RDT-1B. This result is particularly noteworthy given that TwinVLA was pretrained on a substantially smaller single-arm dataset (0.5M vs. RDT-1B’s 1.4M) with far less computational cost, demonstrating the data efficiency of our approach. The comparatively low performance of DP further underscores the critical importance of pretraining. Among all methods, $\pi_{0}$ ultimately demonstrated the best overall performance.
@@ 5.3 Simulation Experiments
RoboTwin 2.0. We use the RoboTwin 2.0 benchmark (chen2025robotwin), consisting of $50$ bimanual tasks. Adhering to the official evaluation protocol, we fine-tune a model per task with $50$ generated demonstrations and perform $100$ test rollouts under both “Easy” and “Hard” settings. For Easy tasks, test scenes match the training data, but the instructions are novel. The Hard tasks introduce variations in texture, object position, and height. For compared methods, we use the results reported from RoboTwin 2.0 (chen2025robotwin).
Tabletop-Sim. To assess dexterous scenarios beyond tasks in RoboTwin, we develop Tabletop-Sim [^1] Our simulation setup is similar to the concurrent work Aloha-Sim, released by Google DeepMind (aloha_sim_github)., a tabletop simulation environment based on dm_control (tunyasuvunakool2020) and assets from ALOHA2 (aloha2team2024aloha2enhancedlowcost) and GSO object dataset (DBLP:journals/corr/abs-2204-11918). We design $5$ representative tasks that require precise bimanual coordination. Specifically, we define four single-tasks and one multi-task: dish-drainer, handover-box, shoes-table, lift-box, and put X box into Y pot. In the “Hard” tasks, we vary background textures and objects. We collect $50$ episodes on each task using absolute EEF control, and fine-tune a model per task, and perform $500$ evaluation rollouts for both “Easy” and “Hard” settings.
@@ 5.4 Data Efficiency
5.5 Language Following Evaluations
A known challenge is that fine-tuning VLMs on robotic data can degrade their ability to faithfully follow nuanced instructions. We therefore evaluate how effectively our model preserves this core capability in a multi-task setting. The “Put X box into Y pot” task involves $3$ box colors (X) and $2$ pot colors (Y), resulting in a total of $6$ possible instruction combinations. As observed in Figure˜8(a), TwinVLA outperforms both RDT-1B and $\pi_{0}$. We believe this performance stems from effectively preserving the knowledge acquired during single-arm pretraining through careful fine-tuning.
@@ 5.6 Policy Robustness
Figure 8: Language following task and ablation results. (a) We evaluate average success rates on the language following tasks in Tabletop-Sim. (b) Ablation studies in the real world and Tabletop-Sim Easy tasks.
@@ 5.7 Ablations
Attention re-weighting. Removing the attention re-weighting mechanism (w/o Re-weighting) increased the initial fine-tuning loss by 40% and decreased final performance by 1.1% and 1.2% in simulation and real world, respectively. This demonstrates that our re-weighting strategy successfully mitigates the input distribution shift between pretraining and fine-tuning.
MoE integration. Building on the previous ablation, we next remove the MoE integration (w/o MoE). This additional change increased the token sequence length by 28% and increased VRAM usage by 21%, making VLA training more burdensome. Surprisingly, it also further decreases the success rate by 1.1% in simulation, suggesting that MoE integration eliminates redundant processing of shared inputs while maintaining the performance.
Joint attention. Lastly, removing the joint attention mechanism (w/o Joint attn) cause the most significant additional performance drops by 4.0% and 8.3% in simulation and real world, respectively. This impact is greater than other components, confirms that the joint attention is the critical mechanism for the bimanual manipulation requires coordination between both arms.
Effect of single-arm pretraining. As a separate, foundational experiment, we assess the role of pretraining by training a model from scratch without OXE dataset (Scratch). This resulted in a 4.6% performance drop in simulation and a stark 32.9% in real world. This result confirms that effective cross-arm coordination is essential for bimanual manipulation and validates joint attention as the critical mechanism for achieving it in our model.
Twin structure. While we have confirmed that joint attention effectively connects the two modules, a crucial question remains: how does this approach compare to a monolithic model that is inherently unified from the start? To answer this, we revisit our comparison against RDT-1B, a monolithic model of a comparable 1.2B parameter size. The results are telling: TwinVLA outperforms RDT-1B by 16.2% in the real world, 5.0% in simulation, and 25.1% in language-following tasks. This provides strong evidence that the inductive bias from the Twin Structure itself is highly beneficial for bimanual manipulation, validating our design choice over a monolithic approach.
@@ 6 Limitations
Since our work adapts a pre-trained single-arm VLA’s representations for bimanual tasks, a current limitation is that the model forgets its single-arm manipulation skills after fine-tuning. Future research into mechanisms that prevent this forgetting could address data scarcity by integrating diverse data, while also improving model explainability and better generalization capability to unseen tasks.
Moreover, the choice of action space is critical for VLA models, particularly for knowledge transfer. We adopt absolute end-effector (EEF) pose control because it provides an embodiment-agnostic representation essential for our single-arm transfer strategy. In contrast, joint positions are specific to a robot’s degrees of freedom (DOF), making them unsuitable for direct policy transfer. We believe that exploring relative absolute actions (Chi-RSS-24) or developing a shared representation across diverse embodiments could enable more efficient transfer.
@@ 7 Conclusion
In this paper, we introduce TwinVLA, a data-efficient VLA model for bimanual manipulation. TwinVLA provides a new perspective on solving bimanual manipulation under scarce bimanual data by leveraging abundant single-arm datasets. From a small amount of bimanual demonstration data, TwinVLA learns to coordinate two copies of a SingleVLA pretrained on large-scale single-arm data via our proposed joint attention. Through exhaustive experiments both in the real world and simulation, TwinVLA demonstrates its data-efficient learning of bimanual tasks compared to prior monolithic approaches. We believe this principle of bridging data availability gaps via leveraging modularity opens up exciting possibilities for other complex robotic domains, such as mobile manipulation, thereby broadening the impact of large-scale robotic learning.
@@ Acknowledgments
This project was supported in part by Microsoft Research Asia and the Microsoft Accelerate Foundation Models Research (AFMR) grant program. This research was also supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants funded by the Korean Government (MSIT) (RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University); RS-2024-00436680, Global Research Support Program in the Digital Field Program), the Alchemist Project (RS-2024-00432143) funded by the Ministry of Trade, Industry & Energy (MOTIE, Korea), and the Electronics and Telecommunications Research Institute (ETRI) grant funded by the Korean government (24ZR1100, A Study of Hyper-Connected Thinking Internet Technology by autonomous connecting, controlling, and evolving ways). LG Electronics provided the Anubis robot, which was used for the experiments.
@@ Appendix A SingleVLA: Efficient Single-Arm Policy Design and Pretraining
To ensure effective transfer to bimanual manipulation, it is crucial to choose an appropriate action space. Heterogeneous joint configurations across robots induce incompatible action spaces and complicate joint training. Prior work mitigates this with robot-specific decoders or high-dimensional zero-padded spaces (nvidia2025gr00tn1openfoundation; doshi2024scalingcrossembodiedlearningpolicy; octo_2023; black2024pi_0; liu2024rdt). Instead, we convert all actions into absolute end-effector (EEF) poses, providing a consistent, semantically meaningful representation across robots that naturally extends to bimanual control. For rotation, we adopt a 6D representation (8953486), which is well suited for neural network learning.
@@ A.1 Pretraining
RT-1 (rt12022arxiv) $24.49\%$
Kuka (filtered) (Yadav_Nagar_Shah_2024) $12.40\%$
BridgeV2 (walke2023bridgedata) $13.74\%$
Taco Play (rosete2022tacorl) $3.10\%$
Jaco Play (dass2023jacoplay) $0.50\%$
Viola (zhu2022viola) $1.00\%$
Berkeley Autolab UR5 (BerkeleyUR5Website) $1.28\%$
Stanford Hydra (belkhale2023hydra) $4.73\%$
Austin Buds (9695333) $0.22\%$
NYU Franka Play (cui2022play) $0.88\%$
FurnitureBench (heo2023furniturebench) $2.40\%$
Austin Sailor (nasiriany2022sailor) $2.33\%$
Austin Sirius (liu2022robot) $1.84\%$
DLR EDAN (shared control) (vogel_edan_2020; quere_shared_2020) $0.05\%$
UT Austin Mutex (shah2023mutex) $2.38\%$
Berkeley FANUC manipulation (zhu2023fanuc) $0.82\%$
CMU Stretch (bahl2023affordances; mendonca2023structured) $0.16\%$
BC-Z (filtered) (jang2021bc) $7.90\%$
FMB (doi:10.1177/02783649241276017) $7.40\%$
Dobb-E (shafiullah2023bringing) $1.50\%$
DROID (khazatsky2024droid) $10.70\%$
SingleVLA is pretrained on an OXE subset (0.5M trajectories); dataset composition and sampling rates appear in Table 1. We adopt the dataset loader from the OpenVLA (kim2024openvla) codebase and apply sampling according to the designated weights. Because some datasets (e.g., Kuka and BC-Z) include failed trajectories, we pre-process to retain only successful ones. All actions are converted to absolute EEF control with 6D rotations, resulting in a 10-Dimensional action space. We further apply frequency matching as described below.
@@
Robotic datasets differ in control frequency, making fixed-length action-chunk prediction misaligned in real time. For example, a 20-step chunk spans $\sim 7$ seconds in RT-1 (rt12022arxiv) (3 Hz) but only $\sim 1.3$ seconds in DROID (khazatsky2024droid) (15 Hz). Mixing low-frequency data like OXE (oxe2024) with high-frequency datasets can degrade pretraining quality. Inspired by $\pi_{0}$-FAST (pertsch2025fastefficientactiontokenization), which uses DCT (1672377) to map $1$-second actions into a consistent space, we perform frequency matching via interpolation: all datasets are resampled to 20 Hz, improving temporal alignment and transfer to high-frequency bimanual tasks.

### v2

@@ 4.3 Mixture-of-Experts Integration
In TwinVLA, feeding shared inputs $(l,I_{\text{ego}})_{t}$ redundantly to both VLMs significantly increases VRAM usage. To address this, we process shared tokens as a single sequence by employing a MoE mechanism that dynamically routes shared tokens between the two VLM experts:
${\color[rgb]{1,0.12109375,0.04296875}\text{MoE}(x)}=w_{\text{left}}\cdot\text{FFN}_{\text{left}}(x)+(1-w_{\text{left}})\cdot\text{FFN}_{\text{right}}(x).$ (3)
For calculating $w_{\text{left}}$, we add a linear layer that takes the embedding as input and outputs the weights via a softmax function. For other components like $\text{Projection},\text{LayerNorm}$, we implement an output-averaging strategy inspired by task arithmetic (Tang et al., 2024). By processing inputs through both backbones and averaging their outputs, we functionally simulate a shared layer without physically merging parameters (see Figure˜3(b) center). This efficient design reduces VRAM usage by 21%, enabling training with a batch size of $8$ on a single $40$ GB GPU.
Attention re-weighting. A potential side effect of introducing new arm-specific tokens is that the model’s learned attention patterns can be disrupted, shifting focus away from the pretrained shared modalities. To mitigate this and preserve the valuable pretrained knowledge, we re-scale the attention scores for the shared modality (Algorithm˜4). This maintains pretrained modality importance, allowing the model to bypass an initial adaptation phase and focus directly on the target task—a benefit evidenced by a lower initial loss and converged loss during fine-tuning.
@@ 5 Experiments
  • Does TwinVLA retain core VLA properties—language-following and robustness to unseen scenes and instructions (Sections˜5.6 and 5.5)?
@@
(a) Real-world tasks
(b) Simulation tasks
Figure 4: Experimental setups. (a) We evaluate TwinVLA on five real-world bimanual tasks using an Anubis robot. (b) We further analyze TwinVLA on a large suite of simulation tasks: 55 tasks in Tabletop-Sim and $50$ tasks in RoboTwin 2.0.
@@ 5.1 Compared Methods
  • RDT-1B (Liu et al., 2024): This serves as our direct baseline. With a comparable size(1.21.2B vs. TwinVLA’s 1.31.3B parameters), it represents the standard monolithic approach that requires substantially larger resources ($\sim$2,400h data, $\sim$11,$440$ H100 days vs. $\sim$800h single-arm data, $\sim$$25$ H100 days).
  • $\bm{\pi_{0}}$ (Black et al., 2024): We include this as an upper-bound, as this is 3.33.3B-parameter VLA trained on over 1010K hours of proprietary robot data. Our goal is to assess how closely TwinVLA can approach this performance ceiling with far greater efficiency.
  • Diffusion Policy (DP) (Chi et al., 2024a): This is a strong baseline method in low-data regime with 271271M parameters, used to demonstrate the crucial benefits of pretraining.
@@ 5.2 Real-World Experiments
Environment. For real-world experiments, we use a dual-arm robot, Anubis (Kang et al., 2025), as shown in Figure˜4(a). Anubis has two $6$ DoF arms with parallel-jaw grippers. The robot is equipped with two wrist-mounted cameras and a single ego-centric view camera.
Tasks. We design five long-horizon tabletop manipulation tasks, which require careful coordination and accurate motions: Fold towel, Extract hexkey, Carrot to bag, Brush to dustpan, and Take towel off, and one task set, Put X into pot. We collect $50$ episodes for each task using absolute EEF control. Each method is fine-tuned for each task and evaluated with $20$ rollouts.
Results. As presented in Figure˜5, our model, TwinVLA, significantly outperforms RDT-1B. This achievement is remarkable considering the data disparity: TwinVLA is pretrained on just $\sim$800h of single-arm data, in contrast to RDT-1B’s usage of a $\sim$2,400h dataset mixed with bimanual trajectories, which highlights the data efficiency of our approach. While DP’s low performance confirms the necessity of pretraining, $\pi_{0}$ achieved the highest overall performance with significantly higher costs.
@@ 5.3 Simulation Experiments
RoboTwin 2.0. We use the RoboTwin 2.0 benchmark (Chen et al., 2025a), consisting of $50$ bimanual tasks. Adhering to the official evaluation protocol, we fine-tune a model per task with $50$ generated demonstrations and perform $100$ test rollouts under both “Easy” and “Hard” settings. For Easy tasks, test scenes match the training data, but the instructions are novel. The Hard tasks introduce variations in texture, object position, and height. For compared methods, we use the results reported from RoboTwin 2.0 (Chen et al., 2025a).
Tabletop-Sim. To assess dexterous scenarios beyond tasks in RoboTwin, we develop Tabletop-Sim [^1] Our simulation setup is similar to the concurrent work Aloha-Sim, released by Google DeepMind (Google DeepMind, 2025)., a tabletop simulation environment based on dm_control (Tunyasuvunakool et al., 2020) and assets from ALOHA2 (Team et al., 2024) and GSO object dataset (Downs et al., 2022). We design $5$ representative tasks that require precise bimanual coordination. Specifically, we define four single-tasks and one multi-task: dish-drainer, handover-box, shoes-table, lift-box, and put X box into Y pot. In the “Hard” tasks, we vary background textures and objects. We collect $50$ episodes on each task using absolute EEF control, and fine-tune a model per task, and perform $500$ evaluation rollouts for both “Easy” and “Hard” settings.
@@ 5.5 Policy Robustness
Table 1: Comparison of success rates for the Fold towel task in challenging scenes.
Model Low light With distractors
RDT 15.0% 15.0%
$\pi_{0}$ 40.0% 60.0%
TwinVLA 45.0% 25.0%
@@
In Table˜1, we additionally compared success rates in unseen real-world settings (see Figure˜13)—specifically low-light and distractor-heavy environments—using the Fold towel task. TwinVLA is robust to lighting changes but less effective with distractors. Meanwhile, $\pi_{0}$ works robustly in both cases, and RDT-1B achieves the lowest success rates.
5.6 Language Following Evaluations
A known challenge is that fine-tuning VLMs on robotic data can degrade their ability to faithfully follow nuanced instructions. We therefore evaluate how effectively our model preserves this core capability in a multi-task setting. We evaluated the “Put X into pot” task across both simulation and real-world settings. As observed in Figure˜8(a), TwinVLA outperforms both RDT-1B and $\pi_{0}$. We believe this performance stems from effectively preserving the knowledge acquired during single-arm pretraining through careful fine-tuning.
@@ 5.6 Language Following Evaluations
Figure 8: Language following task and ablation results. (a) We evaluate average success rates on the language following tasks in the real world and Tabletop-Sim. (b) Ablation studies in the real world and Tabletop-Sim Easy tasks.
@@ 5.7 Ablations
Attention re-weighting. Removing the attention re-weighting mechanism (w/o Re-weighting) increased the initial fine-tuning loss by 40% and decreased final performance by 1.1% and 4.0% in simulation and real world, respectively. This demonstrates that our re-weighting strategy successfully mitigates the input distribution shift between pretraining and fine-tuning.
MoE integration. Building on the previous ablation, we next remove the MoE integration (w/o MoE). This additional change increased the token sequence length by 28% and increased VRAM usage by 21%, making VLA training more burdensome. Surprisingly, it also further decreases the success rate by 1.1% and 5.0%, suggesting that MoE integration eliminates redundant processing of shared inputs while maintaining the performance.
Joint attention. Lastly, removing the joint attention mechanism (w/o Joint attn) causes the most significant additional performance drops of 4.0% and 27.0% in simulation and the real world, respectively. This impact is particularly pronounced in real-world tightly coupled bimanual tasks, confirming that joint attention is a critical mechanism for bimanual coordination.
Effect of single-arm pretraining. As a separate, foundational experiment, we assess the role of pretraining by training a model from scratch without OXE dataset (Scratch). This resulted in a 4.6% performance drop in simulation and a stark 46.0% in real world. This result confirms that effective cross-arm coordination is essential for bimanual manipulation and validates joint attention as the critical mechanism for achieving it in our model.
Twin structure. While we have confirmed that joint attention effectively connects the two modules, a crucial question remains: how does this approach compare to a monolithic model that is inherently unified from the start? To answer this, we revisit our comparison against RDT-1B, a monolithic model of a comparable 1.2B parameter size. The results are telling: TwinVLA outperforms RDT-1B by 26.0% in the real world, 5.0% in simulation, and 21.8% in language-following tasks on average. This provides strong evidence that the inductive bias from the Twin Structure itself is highly beneficial for bimanual manipulation, validating our design choice over a monolithic approach.
@@ 6 Limitations
Generalization remains limited due to the visual disparity of two arms, which differs from the single-arm pretraining distribution. Future research into mechanisms that prevent this could address data scarcity by integrating diverse data, while also improving the model explainability and the better generalization ability to unseen tasks.
Moreover, we adopt absolute end-effector (EEF) pose control, as its embodiment-agnostic nature facilitates single-arm transfer, unlike DOF-specific joint positions. Future exploration of relative absolute actions (Chi et al., 2024b) or shared representations could further enhance transfer efficiency.
@@ 7 Conclusion
In this paper, we introduce TwinVLA, a data-efficient VLA model for bimanual manipulation. TwinVLA provides a new perspective on solving bimanual manipulation under scarce bimanual data by leveraging abundant single-arm datasets. From a small amount of bimanual demonstration data, TwinVLA learns to coordinate two copies of a SingleVLA pretrained on large-scale single-arm data via our proposed method. Through exhaustive experiments both in the real world and simulation, TwinVLA demonstrates its data-efficient learning of bimanual tasks compared to prior monolithic approaches. Beyond the bimanual setting, we believe this work serves as a blueprint for addressing inherent dataset imbalances across modalities. By illustrating how modular relationships can be exploited to bridge these data gaps, TwinVLA opens promising ways for other complex domains—such as mobile manipulation—thereby broadening the impact of large-scale robotic learning.
@@ Acknowledgments
This project was supported in part by Microsoft Research Asia and the Microsoft Accelerate Foundation Models Research (AFMR) grant program. This research was also supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants funded by the Korean Government (MSIT) (RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University); RS-2024-00436680, Global Research Support Program in the Digital Field Program), the Alchemist Project (RS-2024-00432143) funded by the Ministry of Trade, Industry & Energy (MOTIE, Korea), and the Electronics and Telecommunications Research Institute (ETRI) grant funded by the Korean government (26ZR1100, Research on Intelligent Industrial Convergence). LG Electronics provided the Anubis robot, which was used for the experiments. The authors would like to thank Byeongjin Kang for the assistance with the preliminary experiments.
@@ Appendix A SingleVLA: Efficient Single-Arm Policy Design and Pretraining
To ensure effective transfer to bimanual manipulation, it is crucial to choose an appropriate action space. Heterogeneous joint configurations across robots induce incompatible action spaces and complicate joint training. Prior work mitigates this with robot-specific decoders or high-dimensional zero-padded spaces (NVIDIA et al., 2025; Doshi et al., 2024; Octo Model Team et al., 2024; Black et al., 2024; Liu et al., 2024). Instead, we convert all actions into absolute end-effector (EEF) poses, providing a consistent, semantically meaningful representation across robots that naturally extends to bimanual control. For rotation, we adopt a 6D representation (Zhou et al., 2019), which is well suited for neural network learning.
@@ A.1 Pretraining
RT-1 (Brohan et al., 2022) $24.49\%$
Kuka (filtered) (Yadav et al., 2024) $12.40\%$
BridgeV2 (Walke et al., 2023) $13.74\%$
Taco Play (Rosete-Beas et al., 2022) $3.10\%$
Jaco Play (Dass et al., 2023) $0.50\%$
Viola (Zhu et al., 2022a) $1.00\%$
Berkeley Autolab UR5 (Chen et al., 2023) $1.28\%$
Stanford Hydra (Belkhale et al., 2023) $4.73\%$
Austin Buds (Zhu et al., 2022b) $0.22\%$
NYU Franka Play (Cui et al., 2023) $0.88\%$
FurnitureBench (Heo et al., 2023) $2.40\%$
Austin Sailor (Nasiriany et al., 2022) $2.33\%$
Austin Sirius (Liu et al., 2023c) $1.84\%$
DLR EDAN (shared control) (Vogel et al., 2020; Quere et al., 2020) $0.05\%$
UT Austin Mutex (Shah et al., 2023) $2.38\%$
Berkeley FANUC manipulation (Zhu et al., 2023) $0.82\%$
CMU Stretch (Bahl et al., 2023; Mendonca et al., 2023) $0.16\%$
BC-Z (filtered) (Jang et al., 2021) $7.90\%$
FMB (Luo et al., 2025) $7.40\%$
Dobb-E (Shafiullah et al., 2023) $1.50\%$
DROID (Khazatsky et al., 2024) $10.70\%$
SingleVLA is pretrained on an OXE subset ($\sim$800h); dataset composition and sampling rates appear in Table 2. We adopt the dataset loader from the OpenVLA (Kim et al., 2024) codebase and apply sampling according to the designated weights. Because some datasets (e.g., Kuka and BC-Z) include failed trajectories, we pre-process to retain only successful ones. Regarding the action space, we convert all actions to absolute EEF control with 6D rotations. We deliberately selected an absolute representation to mitigate the error accumulation and drift issues often amplified in high-frequency bimanual control. Unlike absolute joint positions, however, absolute EEF poses preserve the embodiment-agnostic property required for heterogeneous pretraining. We define these poses relative to the robot’s base frame, resulting in a 10-Dimensional action space. We further apply frequency matching as described below.
@@
Robotic datasets differ in control frequency, making fixed-length action-chunk prediction misaligned in real time. For example, a 20-step chunk spans $\sim 7$ seconds in RT-1 (Brohan et al., 2022) (3 Hz) but only $\sim 1.3$ seconds in DROID (Khazatsky et al., 2024) (15 Hz). Mixing low-frequency data like OXE (Open X-Embodiment Collaboration et al., 2024) with high-frequency datasets can degrade pretraining quality. Inspired by $\pi_{0}$-FAST (Pertsch et al., 2025), which uses DCT (Ahmed et al., 1974) to map $1$-second actions into a consistent space, we perform frequency matching via interpolation: all datasets are resampled to 20 Hz, improving temporal alignment and transfer to high-frequency bimanual tasks.